OpenStudy (anonymous):

3+√(z-10)=√(z+5) Solve and give solution set.

OpenStudy (tkhunny):

First, write down \(z \ge 10\) and \(z \ge -5\). Why woudl I do that?

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

thats a square root sign

OpenStudy (tkhunny):

Right, so why would I write those things down first, knowing that we have those square roots?

OpenStudy (anonymous):

thats how the problem is written

OpenStudy (tkhunny):

You're not answering my question. I believe how the question is written. It is written quite well. Now answer my question. I wrote \(z \ge -5\) for a very important reason, as part of the solution to this problem. Do you know the reason?

OpenStudy (agent0smith):

Ben, can you take the square root of a negative number? Now think about what tkhunny was telling you. To solve this, it'll help to first square both sides of the equation.